OpenStudy (anonymous):

true or false for any function x=f^-1(y) then x=f^-1(y). explain

OpenStudy (dumbcow):

do you mean "for any function y =f(x) then x = f^-1(y) " ?

OpenStudy (anonymous):

i meant for any function,x=f^-1(y) then y=f(x) explain

OpenStudy (anonymous):

thats what dc was saying. it is true. it is the inverse function; where the x and y values are inversed.

OpenStudy (akshay_budhkar):

do you mean "for any function y =f(x) then x = f^-1(y) " ????

OpenStudy (dumbcow):

It is true because thats how the inverse function is defined: \[x = f(f^{-1}(x))\] If you substitute f^-1(y) for x, then \[y = f(f^{-1}(y))\]